# Augmenting Subspace Optimization Methods with Linear Bandits

### v1

@@ 6 Conclusions and Future Work
Acknowledgment
This work was supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research Applied Mathematics and SciDAC programs Contract Nos. DE-AC02-06CH11357 and DE-AC02-05CH11231.
@@ Appendix A Proof of Lemma 3
$\begin{array}[]{rl}\displaystyle\sum_{j=\ell(k)}^{i}\|s_{j}\|^{2}_{C_{k}^{-1}}&=Tr\left(C_{k}^{-1}\displaystyle\sum_{j=\ell(k)}^{i}s_{j}s_{j}^{\top}\right)\\ &\leq Tr\left(C_{k}^{-1}\displaystyle\sum_{j=\ell(k)}^{i}s_{j}s_{j}^{\top}\right)+\displaystyle\sum_{j=i+1}^{k-1}s_{j}^{\top}C_{k}^{-1}s_{j}+\lambda\displaystyle\sum_{\ell=1}^{d}e_{i}^{\top}C_{k}^{-1}e_{i}\\ &=Tr\left(C_{k}^{-1}\displaystyle\sum_{j=\ell(k)}^{i}s_{j}s_{j}^{\top}\right)+Tr\left(C_{k}^{-1}\displaystyle\sum_{j=1+1}^{k-1}s_{j}s_{j}^{\top}\right)+Tr\left(C_{k}^{-1}\lambda\displaystyle\sum_{\ell=1}^{d}e_{i}e_{i}^{\top}\right)\\ &=Tr(C_{k}^{-1}(A_{k-1}+\lambda I_{d}))=Tr(C_{k}^{-1}C_{k})=Tr(I_{d})=d,\\ \end{array}$
@@
$\begin{array}[]{rl}B_{k}\leq\displaystyle\sum_{i=\ell(k)}^{k-1}\displaystyle\sqrt{\frac{d(i-\ell(k)+1)}{\lambda}}\|\nabla f(x_{i+1})-\nabla f(x_{i})\|&\leq\displaystyle\sqrt{\frac{d(M+1)}{\lambda}}\displaystyle\sum_{i=\ell(k)}^{k-1}\|\nabla f(x_{i+1})-\nabla f(x_{i})\|.\\ \end{array}$
@@ Appendix B Proof of Lemma 4
We begin by constructing a set of auxiliary matrices. Without loss of generality, suppose that $k/M=B$ for some integer $B$. We will consider a partition of $\{1,2,\dots,k\}$ into $B$ consecutive blocks; and for the $b$th of the $B$ blocks and for each $j\in\{1,2,\dots,k\}$, we define the auxiliary matrix

### v4

@@ 6 Conclusions and Future Work
Funding
This work was supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) Programs through Contract Nos. DE-AC02-06CH11357 and through the FASTMath Institute.
Disclosure statement
The authors report there are no competing interests to declare.
Data availability statement
The code used in these experiments will be made available upon reasonable request by contacting the author.
@@ Appendix A Proof of Lemma 3
[…]
@@
[…]
@@ Appendix B Proof of Lemma 4
We begin by constructing a set of auxiliary matrices. Without loss of generality, suppose that $k/M=B$ for some integer $B$. We will consider a partition of $\{1,2,\dots,k\}$ into $B$ consecutive blocks. For the $b$th of the $B$ blocks and for each $j\in\{1,2,\dots,k\}$, we define the auxiliary matrix
@@
The submitted manuscript has been created by UChicago Argonne, LLC, Operator of Argonne National Laboratory (‘Argonne’). Argonne, a U.S. Department of Energy Office of Science laboratory, is operated under Contract No. DE-AC02-06CH11357. The U.S. Government retains for itself, and others acting on its behalf, a paid-up nonexclusive, irrevocable worldwide license in said article to reproduce, prepare derivative works, distribute copies to the public, and perform publicly and display publicly, by or on behalf of the Government. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan. http://energy.gov/downloads/doe-public-access-plan.